# A Fourier-based inference method for learning interaction kernels in particle systems

### v1

@@ 3 Generalized method of moments
and in Algorithm 1 we summarize the main steps needed for its construction. Before studying the dependence of the estimator $(\widehat{W}^{\prime})_{T,N}^{(K)}$ on its parameters $T,N,$ and $K$, in the next section we discuss the problem of inferring the drift term $V^{\prime}$.
@@ 3.2 Convergence analysis for the interaction kernel estimator
    In this section, we study the convergence properties of the estimator $(\widehat{W}^{\prime})_{T,N}^{(K)}$ defined in equation (3.13) as $T,N,K\to\infty$. We note that $K$ is the number of Fourier coefficients that we use to approximate the interaction kernel. Therefore, fixing $K$ gives the best possible approximation that can be reached in the ideal setting in which we observe an infinite trajectory from a system of infinite particles. On the other hand, we will observe that increasing $K$ results in a worse conditioning of the linear system to be solved, which in turn implies that larger values of $T$ and $N$ are necessary in order to obtain an accurate approximation of the interaction kernel. Our study follows the approach in [40], and is based on a forward error stability analysis for the linear system (3.10), which can be seen as a perturbation of equation (3.8). In the next lemma, whose proof is inspired by the proof of [40, Lemma 4.5], we quantify the error that we commit in replacing the exact moments with the empirical moments on both the left-hand side and the right-hand side of the linear system.
@@
    Under Assumptions 1.1, 2.4, and 3.1, for all $K\geq 0$ there exists a constant $C=C(K)>0$, independent of $T$ and $N$, such that
@@
    $\operatorname{\mathbb{E}}\left[\left\|\widetilde{\gamma}^{(K)}_{T,N}-\gamma^{(K)}\right\|\right]$ $\leq\sum_{i=0}^{K}\sum_{j=1}^{i}j\left(\operatorname{\mathbb{E}}\left[\frac{1}{\widetilde{c}_{i}}\left\lvert\widetilde{\lambda}_{ij}-\lambda_{ij}\right\rvert\left\lvert\widetilde{\mathbb{M}}^{(j-1)}\right\rvert\right]+\left\lvert\lambda_{ij}\right\rvert\operatorname{\mathbb{E}}\left[\frac{1}{\widetilde{c}_{i}}\left\lvert\widetilde{\mathbb{M}}^{(j-1)}-\mathbb{M}^{(j-1)}\right\rvert\right]\right)$ (3.22)
    $\quad+\sum_{i=0}^{K}\operatorname{\mathbb{E}}\left[\left\lvert\frac{1}{\widetilde{c}_{i}}-\frac{1}{c_{i}}\right\rvert\right]\left(\sum_{j=1}^{i}j\left\lvert\lambda_{ij}\mathbb{M}^{(j-1)}\right\rvert\right).$
    Then, using Hölder’s inequality we obtain for $r$ given by Assumption 2.4
@@
    ×$\qquad\times\left(\operatorname{\mathbb{E}}\left[\left\lvert\widetilde{\mathbb{M}}^{(j-1)}\right\rvert^{\frac{r(3r-2)}{(r-2)(r-1)}}\right]\right)^{\frac{(r-2)(r-1)}{r(3r-2)}},$
@@
    Therefore, using bounds (3.23) and (3.24) in equation (3.22), due to the boundedness of the moments, equation (2.8), Assumption 2.4, and Lemmas 2.2 and 2.6, we deduce (i). We proceed similarly for (ii) and we get
@@
    $\quad+\sum_{i=0}^{K}\operatorname{\mathbb{E}}\left[\left\lvert\frac{1}{\widetilde{c}_{i}}-\frac{1}{c_{i}}\right\rvert\right]\left(\sum_{j=0}^{i}\left\lvert\lambda_{ij}\operatorname{\mathbb{E}}^{\rho}\left[V^{\prime}(X)X^{j}\right]\right\rvert\right).$
@@
    and the bound (3.27) follows since $V^{\prime}$ is polynomially bounded and due to the boundedness of the moments. For equation (3.28), we proceed as in the proof of [40, Lemma 4.3] and we obtain
@@
    Then, for $E_{1}$ and $E_{3}$, still following the proof of [40, Lemma 4.3], we apply the mean ergodic theorem in [33, Section 4], Hölder’s and Jensen’s inequality, the boundedness of the moments, and the uniform propagation of chaos property to get $E_{1}\leq\widetilde{C}/\sqrt{T}$ and $E_{3}\leq\widetilde{C}/\sqrt{N}$. Regarding $E_{2}$, using the mean value theorem, we have
@@
    $\operatorname{\mathbb{E}}\left[\left\|\widetilde{B}^{(K)}_{T,N}-B^{(K)}\right\|\right]$ $\leq\sum_{i=0}^{K}\sum_{k=0}^{K}\operatorname{\mathbb{E}}\left[\left|\frac{1}{\widetilde{c}_{i}\widetilde{c}_{k}}\sum_{\ell=0}^{i}\sum_{j=0}^{k}\sum_{m=0}^{j}(-1)^{j-m}\binom{j}{m}\widetilde{\lambda}_{i\ell}\widetilde{\lambda}_{kj}\widetilde{\mathbb{M}}^{(m+\ell)}\widetilde{\mathbb{M}}^{(j-m)}\right.\right.$ (3.32)
@@
    respectively, and then using the boundedness of the moments, equation (2.8), Assumption 2.4, and Lemmas 2.2 and 2.6, we obtain (iii), which concludes the proof. ∎
@@
    Under Assumptions 1.1 and 3.1, it holds
@@
    We can now estimate the error of the estimator $\widehat{\beta}_{T,N}^{(K)}$ in equation (3.12) with respect to the true Fourier coefficients $\beta^{(K)}$. The proof of the next result is based on [42, Section 3.1.2] and [40, Theorem 4.6].
@@
    Under Assumptions 1.1, 2.4, and 3.1, for all $K\geq 0$, if $\det(B^{(K)})\neq 0$, then there exists a constant $C=C(K)>0$, independent of $T$ and $N$, such that
@@
    and due to Markov’s inequality and Lemma 3.5(iii) we have
@@
    ×$\qquad\times\left[(B^{(K)})^{-1}\left(\sigma(\widetilde{\gamma}_{T,N}^{(K)}-\gamma^{(K)})-(\widetilde{\alpha}_{T,N}^{(K)}-\alpha^{(K)})\right)+(B^{(K)})^{-1}e^{(K)}\right],$
    which, following the proof of [42, Theorem 3.1], implies
    $\operatorname{\mathbb{E}}\left[\left\|\widetilde{\beta}_{T,N}^{(K)}-\beta^{(K)}\right\||A_{K}\right]$ $\leq 2\left\|(B^{(K)})^{-1}\right\|\operatorname{\mathbb{E}}\left[\sigma\left\|\gamma_{T,N}^{(K)}-\gamma^{(K)}\right\|+\left\|\alpha_{T,N}^{(K)}-\alpha^{(K)}\right\||A_{K}\right]$ (3.47)
    $\quad+2\left\|(B^{(K)})^{-1}e^{(K)}\right\|.$
    Using the fact that $\operatorname{\mathbb{E}}[Z|A_{K}]\leq\operatorname{\mathbb{E}}[Z]/\mathbb{P}(A_{K})$ for a positive random variable $Z$, applying Lemma 3.5, and due to equation (3.44), we get for $T$ and $N$ sufficiently large
@@
    Finally, in the next result we consider the estimator $(\widehat{W}^{\prime})_{T,N}^{(K)}$ in equation (3.13), and, employing Proposition 3.7, we analyze the approximation error with respect to the true interaction kernel $W^{\prime}$.
@@
    Under Assumptions 1.1, 2.4, and 3.1, for all $K\geq 0$, if $\det(B^{(K)})\neq 0$, then there exists a constant $C=C(K)>0$, independent of $T$ and $N$, such that
@@
    which, due to Propositions 2.7 and 3.7, gives the desired result. ∎
@@
    In order for the estimator to converge, the additional terms $\delta(K)$ and $\epsilon(K)$ in Theorem 3.8 must disappear as $K$ increases. First, notice that $\epsilon(K)\to 0$ as $K\to\infty$ because $W^{\prime}\in L^{2}(\rho)$. In principle, we should be able to relate the decay of the generalized Fourier coefficients of the interaction kernel (or its derivative) with the regularity of $W$, using results from approximation theory such as Jackson’s inequality, as was done in [21]. In our setting, we would need to study this problem in weighted Sobolev spaces on the whole real line, but the study of this interesting question is beyond the scope of this paper. Regarding $\delta(K)$, since, by Lemma 3.6, $\|e^{(K)}\|$ vanishes, it suffices to require that $\|(B^{(K)})^{-1}\|$ do not blow up faster than the rate at which $\|e^{(K)}\|$ converges to zero. The proof of this fact is nontrivial, as it depends on the unknown interaction kernel $W^{\prime}$. However, we observed that this was indeed the case in all of the numerical experiments that we considered. In particular, this is straightforward to verify for all polynomial interactions, since $e^{(K)}=0$ (and also $\epsilon(K)=0$) for all $K\geq r$ with $r$ being the degree of the polynomial, which in turn implies $\delta(K)=0$. Alternatively, notice that we can write
@@ 4 Numerical experiments
    In this section, we employ the methodology introduced above to infer interaction kernels in particle systems, and verify numerically the estimates predicted by the theory. We first consider the mean-field Ornstein–Uhlenbeck process, for which a basis of orthogonal polynomials with respect to the invariant measure can be computed analytically. Even though the Ornstein–Uhlenbeck process is a simple test case, it is still an interesting example because it allows us to assess the performance of our method, since both the invariant measure and the corresponding orthogonal polynomials are given in closed form. Then, we consider more complex interaction kernels, including examples where the assumptions of our theoretical analysis are not necessarily satisfied. Even in this case, numerical experiments demonstrate that our methodology can still be used to learn the interaction kernel from a single trajectory. We generate synthetic observations by numerically solving the interacting particle system (1.1) with deterministic initial conditions, $X_{0}^{(n)}=0$ for all $n=1,\dots,N$. The SDE system is discretized employing the Euler–Maruyama scheme with a time step $h=0.01$. Then, we assume to observe only the first particle in the system to infer the interaction kernel, so that $Y_{t}=X_{t}^{(1)}$ for all $t\in[0,T]$.
@@ 4.1 Mean-field Ornstein–Uhlenbeck process
[…]
    Figure 2: Comparison between the theoretical and empirical rate of convergence of the first four (excluding the constant function) orthogonal polynomials with respect to the invariant measure $\rho$ of the mean-field Ornstein–Uhlenbeck process in $L^{2}(\rho)$, for both the number of particles $N$ (left) and the final time $T$ (right).
[…]
@@
[…]

### v2

@@ 3 Generalized method of moments
Remark 3.3.
To solve the linear system (3.10), the diffusion coefficient $\sigma$ is assumed to be known, since this work focuses on the inference problem for the interaction kernel. Nevertheless, it is also possible to identify both the diffusion coefficient and the interaction kernel sequentially. Specifically, given a particle trajectory $(Y_{t})_{t\in[0,T]}$ as defined in equation (1.7), the diffusion coefficient can be estimated using the quadratic variation (see, e.g., [42, Section 5.3])
$\mathcal{Q}^{\Delta}_{T,N}=\sum_{j=1}^{J}(Y_{j\Delta}-Y_{(j-1)\Delta})^{2},$ (3.12)
where $\Delta>0$ is the time step and $J=T/\Delta$. In fact, it holds that
$\lim_{\Delta\to 0}\mathcal{Q}^{\Delta}_{T,N}=2\sigma T,\qquad\text{a.s.},$ (3.13)
independently of $N$, which implies that an estimator for the diffusion coefficient is given by
$\widehat{\sigma}^{\Delta}_{T,N}=\frac{\mathcal{Q}^{\Delta}_{T,N}}{2T}.$ (3.14)
In addition, our spectral estimator from [44] can be used to infer the diffusion coefficient using low-frequency observations. The estimated diffusion coefficient can then be used in the subsequent estimation of the interaction kernel. A natural direction for future research is to rigorously analyze how the error in estimating the diffusion coefficient propagates to the kernel estimator. We expect this to lead to an additional error term in the convergence analysis, which would appear on the right-hand side of Theorem˜3.10.
@@
and in Algorithm˜1 we summarize the main steps needed for its construction. We emphasize again that the system size $N$ is not required to infer the interaction kernel. The subscript $N$ in the estimator $(\widehat{W}^{\prime})_{T,N}^{(K)}$ simply indicates that it is based on observations generated by an interacting particle system with $N$ particles, and therefore it depends on $N$ only implicitly.
@@
  • Empirical moments $\widetilde{\mathbb{M}}_{T,N}^{(r)}$Orthonormal basis $\widetilde{\psi}_{k}$Linear system $\widehat{\beta}^{(K)}_{T,N}$Interaction estimator $(\widehat{W}^{\prime})_{T,N}^{(K)}$
@@
    Before studying the dependence of the estimator $(\widehat{W}^{\prime})_{T,N}^{(K)}$ on its parameters $T,N,$ and $K$, in the next section we discuss the problem of inferring the drift term $V^{\prime}$.
@@ 3.1 Inference of the drift term
    Remark 3.6.
    In principle, our methodology could recover both $V^{\prime}$ and $W^{\prime}$ if one further assumes that the subspaces spanned by their generalized Fourier coefficients are orthogonal, i.e., that the sets of non-zero Fourier coefficients with respect to the orthonormal basis are disjoint. However, this is not a realistic assumption in most practical settings
@@ 3.2 Convergence analysis for the interaction kernel estimator
    In this section, we study the convergence properties of the estimator $(\widehat{W}^{\prime})_{T,N}^{(K)}$ defined in equation (3.16) as $T,N,K\to\infty$. We note that $K$ is the number of Fourier coefficients that we use to approximate the interaction kernel. Therefore, fixing $K$ gives the best possible approximation that can be reached in the ideal setting in which we observe an infinite trajectory from a system of infinite particles. On the other hand, we will observe that increasing $K$ results in a worse conditioning of the linear system to be solved, which in turn implies that larger values of $T$ and $N$ are necessary in order to obtain an accurate approximation of the interaction kernel. Our study follows the approach in [45], and is based on a forward error stability analysis for the linear system (3.10), which can be seen as a perturbation of equation (3.8). In the next lemma, whose proof is inspired by the proof of [45, Lemma 4.5], we quantify the error that we commit in replacing the exact moments with the empirical moments on both the left-hand side and the right-hand side of the linear system.
@@
    Let $\gamma^{(K)}$, $\alpha^{(K)}$, $B^{(K)}$ be defined in equations (3.6), (3.7), (3.8), and let $\widetilde{\gamma}^{(K)}_{T,N}$, $\widetilde{\alpha}^{(K)}_{T,N}$, $\widetilde{B}^{(K)}_{T,N}$ be their approximations defined in equation (3.11). Under Assumptions˜1.1, 2.4, and 3.1, for all $K\geq 0$ there exists a constant $C=C(K)>0$, independent of $T$ and $N$, such that
@@
[…]
    Then, using Hölder’s inequality we obtain for $r$ given by Assumption˜2.4
@@
    $\qquad\times\left(\operatorname{\mathbb{E}}\left[\left\lvert\widetilde{\mathbb{M}}^{(j-1)}\right\rvert^{\frac{r(3r-2)}{(r-2)(r-1)}}\right]\right)^{\frac{(r-2)(r-1)}{r(3r-2)}},$
@@
    Therefore, using bounds (3.26) and (3.27) in equation (3.25), due to the boundedness of the moments, equation (2.8), Assumption˜2.4, and Lemmas˜2.2 and 2.6, we deduce (i). We proceed similarly for (ii) and we get
@@
[…]
@@
    and the bound (3.30) follows since $V^{\prime}$ is polynomially bounded and due to the boundedness of the moments. For equation (3.31), we proceed as in the proof of [45, Lemma 4.3] and we obtain
@@
    Then, for $E_{1}$ and $E_{3}$, still following the proof of [45, Lemma 4.3], we apply the mean ergodic theorem in [37, Section 4], Hölder’s and Jensen’s inequality, the boundedness of the moments, and the uniform propagation of chaos property to get $E_{1}\leq\widetilde{C}/\sqrt{T}$ and $E_{3}\leq\widetilde{C}/\sqrt{N}$. Regarding $E_{2}$, using the mean value theorem, we have
@@
[…]
@@
[…]
    respectively, and then using the boundedness of the moments, equation (2.8), Assumption˜2.4, and Lemmas˜2.2 and 2.6, we obtain (iii), which concludes the proof. ∎
@@
    Let $e^{(K)}$ be defined in equation (3.9). Under Assumptions˜1.1 and 3.1, it holds
@@
    We can now estimate the error of the estimator $\widehat{\beta}_{T,N}^{(K)}$ in equation (3.15) with respect to the true Fourier coefficients $\beta^{(K)}$. The proof of the next result is based on [47, Section 3.1.2] and [45, Theorem 4.6].
@@
    Let $\beta^{(K)}$ be defined in equation (3.9), and let $\widehat{\beta}_{T,N}^{(K)}$ be the estimator given in equation (3.15). Under Assumptions˜1.1, 2.4, and 3.1, for all $K\geq 0$, if $\det(B^{(K)})\neq 0$ where $B^{(K)}$ is defined in equations (3.7), (3.8), then there exists a constant $C=C(K)>0$, independent of $T$ and $N$, such that
@@
    and due to Markov’s inequality and Lemma˜3.7(iii) we have
@@
    $\qquad\times\left[(B^{(K)})^{-1}\left(\sigma(\widetilde{\gamma}_{T,N}^{(K)}-\gamma^{(K)})-(\widetilde{\alpha}_{T,N}^{(K)}-\alpha^{(K)})\right)+(B^{(K)})^{-1}e^{(K)}\right]$
    $\quad-\left(\mathcal{I}_{K+1}+(B^{(K)})^{-1}(\widetilde{B}_{T,N}^{(K)}-B^{(K)})\right)^{-1}(B^{(K)})^{-1}(\widetilde{B}_{T,N}^{(K)}-B^{(K)})\beta^{(K)},$
    which, following the proof of [47, Theorem 3.1], implies
    $\operatorname{\mathbb{E}}\left[\left\|\widetilde{\beta}_{T,N}^{(K)}-\beta^{(K)}\right\||A_{K}\right]$ $\leq 2\left\|(B^{(K)})^{-1}\right\|\operatorname{\mathbb{E}}\left[\sigma\left\|\widetilde{\gamma}_{T,N}^{(K)}-\gamma^{(K)}\right\|+\left\|\widetilde{\alpha}_{T,N}^{(K)}-\alpha^{(K)}\right\||A_{K}\right]$ (3.50)
    $\hskip-5.69046pt+2\left\|(B^{(K)})^{-1}e^{(K)}\right\|+2\left\|(B^{(K)})^{-1}\right\|\left\|\beta^{(K)}\right\|\operatorname{\mathbb{E}}\left[\left\|\widetilde{B}_{T,N}^{(K)}-B^{(K)}\right\||A_{K}\right].$
    Using the fact that $\operatorname{\mathbb{E}}[Z|A_{K}]\leq\operatorname{\mathbb{E}}[Z]/\mathbb{P}(A_{K})$ for a positive random variable $Z$, applying Lemma˜3.7, and due to equation (3.47), we get for $T$ and $N$ sufficiently large
@@
    Finally, in the next result we consider the estimator $(\widehat{W}^{\prime})_{T,N}^{(K)}$ in equation (3.16), and, employing Proposition˜3.9, we analyze the approximation error with respect to the true interaction kernel $W^{\prime}$.
@@
    Let $(\widehat{W}^{\prime})_{T,N}^{(K)}$ be the interaction kernel estimator defined in equation (3.16). Under Assumptions˜1.1, 2.4, and 3.1, for all $K\geq 0$, if $\det(B^{(K)})\neq 0$ where $B^{(K)}$ is defined in equations (3.7), (3.8), then there exists a constant $C=C(K)>0$, independent of $T$ and $N$, such that
@@
    which, due to Propositions˜2.7 and 3.9, gives the desired result. ∎
@@
    In order for the estimator to converge, the additional terms $\delta(K)$ and $\epsilon(K)$ in Theorem˜3.10 must disappear as $K$ increases. First, notice that $\epsilon(K)\to 0$ as $K\to\infty$ because $W^{\prime}\in L^{2}(\rho)$. In principle, we should be able to relate the decay of the generalized Fourier coefficients of the interaction kernel (or its derivative) with the regularity of $W$, using results from approximation theory such as Jackson’s inequality, as was done in [22]. In our setting, we would need to study this problem in weighted Sobolev spaces on the whole real line, but the study of this interesting question is beyond the scope of this paper. Regarding $\delta(K)$, since, by Lemma˜3.8, $\|e^{(K)}\|$ vanishes, it suffices to require that $\|(B^{(K)})^{-1}\|$ do not blow up faster than the rate at which $\|e^{(K)}\|$ converges to zero. The proof of this fact is nontrivial, as it depends on the unknown interaction kernel $W^{\prime}$. However, we observed that this was indeed the case in all of the numerical experiments that we considered. In particular, this is straightforward to verify for all polynomial interactions, since $e^{(K)}=0$ (and also $\epsilon(K)=0$) for all $K\geq r$ with $r$ being the degree of the polynomial, which in turn implies $\delta(K)=0$. Alternatively, notice that we can write
@@ 4 Numerical experiments
    In this section, we employ the methodology introduced above to infer interaction kernels in particle systems, and verify numerically the estimates predicted by the theory. We first consider the mean-field Ornstein–Uhlenbeck process, for which a basis of orthogonal polynomials with respect to the invariant measure can be computed analytically. Even though the Ornstein–Uhlenbeck process is a simple test case, it is still an interesting example because it allows us to assess the performance of our method, since both the invariant measure and the corresponding orthogonal polynomials are given in closed form. Then, we consider more complex interaction kernels, including examples where the assumptions of our theoretical analysis are not necessarily satisfied. Even in this case, numerical experiments demonstrate that our methodology can still be used to learn the interaction kernel from a single trajectory. We generate synthetic observations by numerically solving the interacting particle system (1.1) with deterministic initial conditions, $X_{0}^{(n)}=0$ for all $n=1,\dots,N$. The SDE system is discretized employing the Euler–Maruyama scheme with a time step $h=0.01$. Then, we assume to observe only the first particle in the system to infer the interaction kernel, so that $Y_{t}=X_{t}^{(1)}$ for all $t\in[0,T]$. We notice that, while $N$ is used in the simulation to generate the trajectories, it is not needed for the estimation once the observations are available. In fact, our approach assumes access to a single observed trajectory, with no additional information about the system size.
@@ 4.1 Mean-field Ornstein–Uhlenbeck process
    Figure 2: Comparison between the first four (excluding the constant function) exact ($\psi_{k}$) and approximated ($\widetilde{\psi}_{k}$) orthogonal polynomials with respect to the invariant measure $\rho$ of the mean-field Ornstein–Uhlenbeck process. Top: we fix T=10  [^0] T=10\,000 and vary $N=5,50,500$. Bottom: we fix $N=500$ and vary T=100,1 000,10  [^0] T=100,1\,000,10\,000.
    Figure 3: Comparison between the theoretical and empirical rate of convergence of the first four (excluding the constant function) orthogonal polynomials with respect to the invariant measure $\rho$ of the mean-field Ornstein–Uhlenbeck process in $L^{2}(\rho)$, for both the number of particles NN (left) and the final time $T$ (right).
    Figure 4: Comparison between the true interaction kernel $W^{\prime}$ and the estimators $(\widehat{W}^{\prime})_{T,N}^{(K)}$ in two different cases (T=1 000,N=50T=1\,000,N=50 and T=10 000,N=500T=10\,000,N=500), for different numbers of Fourier coefficients $K=1,\dots,8$ for the Ornstein-Uhlenbeck interaction kernel.
@@
    In Figures˜2 and 3 we compare the exact orthogonal polynomials of equation (4.4) with the approximated polynomials obtained using the single-particle trajectory observation, for different values of the number of particles $N$ in the system and the final observation time $T$. In particular, in Figure˜2 we plot the first four polynomials (starting at $k=1$) varying $N=5,50,500$ with T=10  [^0] T=10\,000 fixed and then varying T=100,1 000,10  [^0] T=100,1\,000,10\,000 with $N=500$ fixed. We observe that, as expected, the approximation error improves when $T$ and $N$ are larger. Moreover, in Figure˜3 we compute the approximation error in the space $L^{2}(\rho)$ and verify the convergence rate provided by the theory. We remark that, even if the error is computed for a single observation and the theoretical rate holds in expectation, the two rates match. We also notice that the approximation error is greater for polynomials with higher degree $k$, and this is due to the constant $C(k)$ in Proposition˜2.7 which grows for larger values of $k$.
    We then apply the proposed methodology to learn the interaction kernel $W^{\prime}(x)=x$, and we consider two cases: “few” observations (T=1  [^0] T=1\,000) in a small system ($N=50$) and “many” observations (T=10  [^0] T=10\,000) in a large system ($N=500$). In Figure˜4 we compare the results for different Fourier series truncations $K=1,\dots,8$ used in the expansion of the interaction kernel. We note that, in this simple setting, two Fourier coefficients ($\beta_{0}$ and $\beta_{1}$) are enough to approximate $W^{\prime}$ and $\delta(K)=\epsilon(K)=0$ in the statement of Theorem˜3.10, since $\beta_{k}=0$ for all $k\geq 2$. Therefore, we notice that increasing $K$ only worsens the results due to the constant $C(K)$, appearing in front of $1/\sqrt{T}$ and $1/\sqrt{N}$, which blows up for larger values of $K$. This is due to the ill-conditioning of both the linear system used to compute the estimator, specifically the fact that the condition number of the matrix $B^{(K)}$ increases with $K$, and the Gram–Schmidt algorithm applied to obtain the orthonormal basis. On the other hand, we observe that if the time of observation and the number of particles in the system are larger, then polynomials of higher degree still provide accurate approximations of the interaction kernel, and this is in agreement with the theoretical result in Theorem˜3.10. We emphasize that this numerical experiment shows the importance of the choice of $K$ in the method, in case we have limited observations and/or small interacting particle systems, since numerical instability becomes immediately apparent in such cases. Therefore, it would be interesting to determine the criteria for automatically adjusting $K$, and we will return to this problem in future work.
    4.2 Discrete-time, low-frequency observations
    Figure 5: Comparison between the true interaction kernel $W_{0}^{\prime}$ from equation (4.5) and the estimator $(\widehat{W}_{0}^{\prime})_{T,N}^{(K)}$, for the case of discrete-time observations with different sampling rates $\Delta=1,2,4,8$.
    As highlighted in Remark˜2.1, the methodology developed in this work can also be applied when only discrete-time observations are available. In this section, we present a numerical test case to illustrate this setting. We consider a system with diffusion coefficient $\sigma=1$, a quadratic confining potential $V(x)=x^{2}/2$, and a polynomial interaction potential with sinusoidal term given by
    $W_{0}(x)=\frac{x^{4}}{4}-\frac{x^{3}}{3}+\frac{x^{2}}{2}+10\cos(x).$ (4.5)
    We set the final time to $T=5000$, the number of particles to $N=250$, and the number of Fourier coefficients to $K=4$.
    We assume that only discrete-time samples $\{\widetilde{Y}_{i}\}_{i=0}^{I}$ from a single particle trajectory of the interacting particle system (as defined in equation (1.7)) are available, for different sampling intervals $\Delta=1,2,4,8$. The corresponding estimators $(\widehat{W}^{\prime}_{i})_{T,N}^{(K)}$ are shown in Figure˜5, where they are compared against the true interaction kernel. We observe that the estimator accurately recovers the interaction kernel for all tested values of $\Delta$. In particular, the quality of the reconstruction does not appear to degrade significantly with larger sampling intervals, indicating robustness of the method to low-frequency data.